Please help me solve this word problem.
Suppose that coffee that is 155°F is poured into a coffee mug and allowed to cool in a room where the temperature is 65°F. The temperature T(in°F) of the coffee at time t ( in minutes) can be modeled by T(t) = 65 + 90e - 0.045t
According to the model when will the temperature of the coffee be 140°F?

Answer by Boreal(15235)

140=65+90e^(-0.045t)
75=90e^(-0.045t)
(75/90)=e^-0.045t=0.83333
ln both sides to remove e
-0.1823=-0.045 t
t=4.05 minutes or about 4m3s